Izzy bought a house for $200,000 which she financed for 30 years at 3%, annual payments. How much Principal did she pay on the loan the first year?

$4,203.85

$6,000.00

$8,319.61

$10,203.85
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

200,000=Annuity[1-(1.03)^-30]/0.03

200,000=Annuity*19.6004413

Annuity=200,000/19.6004413

=$10203.85(Approx)

Interest payment for the first year=$200,000*3%=$6000

Hence principal payment for the first year=Total payment-Interest payment

=(10203.85-6000)

=$4203.85(Approx).